A random variable X has pdf:

X


1

2

3

f(x)

1/8

3/8

3/8

1/8



1
Expert's answer
2021-09-16T00:46:22-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & 0 & 1 & 2 & 3 \\ \hline
 f(x) & 1/8 & 3/8 & 3/8 & 1/8 \\
 
\end{array}"

a)

"F(0)=P(X\leq 0)=P(X=0)"

"=\dfrac{1}{8}"

"F(1)=P(X\leq 1)=P(X=0)+P(X=1)"

"=\dfrac{1}{8}+\dfrac{3}{8}=\dfrac{1}{2}"

"F(2)=P(X\leq2)=P(X=0)+P(X=1)+P(X=2)"

"=\dfrac{1}{8}+\dfrac{3}{8}+\dfrac{3}{8}=\dfrac{7}{8}"

"F(1)=P(X\leq 3)=P(X=0)+P(X=1)"

"+P(X=2)+P(X=3)=\dfrac{1}{8}+\dfrac{3}{8}+\dfrac{3}{8}+\dfrac{1}{8}=1"

"F(x) = \begin{cases}
 0 &x<0 \\
 1/8 & 0\leq x<1\\
 1/2 & 1\leq x<2\\
 7/8 & 2\leq x<3\\
 1 & x\geq 3\\
\end{cases}"

b)